OpenStudy (anonymous):

Solve the equation (4x + 1)(3x - 2) = 0 (Points : 1) A: {-4, 1.5} B: {-1/4, 2/3} C: {1/4, 2/3} D: {-4, 3}

OpenStudy (anonymous):

Same method!!!!

OpenStudy (anonymous):

4x + 1=0 Solve for x

OpenStudy (angela210793):

or 3x-2=0

OpenStudy (anonymous):

-1/4

OpenStudy (anonymous):

Yeah but that solution is not given as one of the answers

OpenStudy (anonymous):

right? for x is - 1/4

OpenStudy (anonymous):

Great johny find the point where x=-1/4

OpenStudy (anonymous):

Got it romero Thanks for the help

OpenStudy (anonymous):

It's B!

OpenStudy (anonymous):

yeah I found it already lol

OpenStudy (anonymous):

You can always check your answer by plugging x=-1/4 back to the original equation. If you get the same Y ( 2/3) Then you got it right,

OpenStudy (anonymous):

I was just answering a message someone sent me

OpenStudy (anonymous):

That's to double check. Doesn't hurt to try. You never want to be wrong :/

OpenStudy (anonymous):

Thanks

OpenStudy (anonymous):

so 4 times -1/4 is -1 +1 is zero so it fits